# Diffusion-based Task-oriented Semantic Communications with Model Inversion Attack

### v1

@@ IV-C Results on CIFAR10 dataset
We also implement the system on the CIFAR10 dataset, which has complex scenes that pose more challenges for semantic communication systems. Fig. 9 illustrates the receiver’s classification accuracy and the attacker’s success rate as functions of $t$. Given our assumption that the training of the $\mathcal{T}-\mathcal{R}$ pair takes precedence over the adversary’s training, even at $t=0$, the classification accuracy of the adversary’s reconstructed images is not very high. This is because the $\mathcal{T}-\mathcal{R}$ pair’s training prioritizes the completion of the label classification task, and the backpropagation process optimizes the $\mathcal{T}$’s neural network in a way that is not aligned with the adversary’s image recovery requirements, which is acceptable since normal systems are not designed to meet attackers’ needs.
@@
The two red lines in Fig. 9 show that the system incorporating the diffusion module outperforms the classifier-only scenario. This indicates that imperfect labels can still provide coarse-grained constraints for the diffusion path. However, the additional performance improvement brought by E-DiffSem is less pronounced compared to the MNIST dataset. This discrepancy may be because the CIFAR10 dataset contains richer and more complex information than MNIST, which makes it more challenging for the imperfect labels estimated by the classifier to provide effective guidance to the diffusion module. In the future, we could explore integrating Transformer-based self-attention mechanisms or leveraging image semantic segmentation techniques to provide stronger guidance for the diffusion module, thereby further enhancing system performance.
@@
Figure 10: Plot of accuracy versus SSIM and PSNR using CIFAR10 dataset.
Figure 11: Adversary’s output images with different $t$ under CIFAR10 dataset. (a) Original image. (b)(c)(d) Adversary’s recovery when $t=60$, $150$ and $300$.
Fig. 10 shows the accuracy versus SSIM and PSNR on the CIFAR10 dataset. It highlights the weak correlation between traditional image quality metrics and semantic security. For instance, when SSIM is $0.12$, the attacker’s category accuracy still reaches $51.8\%$, whereas in the direct mode, with SSIM equal to $0.14$, the accuracy is only $42.6\%$. This difference arises from the attacker’s ability to implicitly learn class-discriminative features, while variations in pixel texture are insufficient to prevent the attacker from acquiring these features. Other curves in this figure can further demonstrate that SSIM and PSNR are not suitable metrics for evaluating the amount of task-relevant information obtained by the attacker in task-oriented communication systems.
As a supplement to the discussion, here we remark the computational complexity of our proposed system. Our hardware conditions are: CPU, Intel Xeon Gold 5220; GPU, GeForce RTX 3090. Although the diffusion model introduces additional computational overhead, the DDIM acceleration strategy restricts the additional single-sample inference time of DiffSem to approximately $0.1$s per batch (batch size 256) on CIFAR10. It shows that the performance enhancement does not impose a substantial computational burden given the relatively small dataset. For higher-resolution datasets, such as ImageNet, the trade-off between performance improvement and computational cost may become a critical consideration, especially in real-time applications where latency is a key factor. Future work could focus on optimizing the diffusion process, potentially exploring techniques such as knowledge distillation, model pruning, or adopting more efficient neural network architectures to address this challenge.
@@ V Conclusion
In this paper, we have introduced a diffusion-based semantic communication framework named DiffSem, to address the dual challenges of optimizing task performance and ensuring privacy protection in semantic communication systems. By leveraging the diffusion mechanism and integrating label embedding, DiffSem effectively reconstructs semantic information and significantly improves task accuracy, as evidenced by a $10.3\%$ boost on the MNIST dataset. To tackle the growing concern of privacy risks posed by model inversion attacks, we proposed a novel evaluation metric that quantifies the semantic fidelity of adversarial outputs, offering a precise, task-relevant method to assess privacy vulnerabilities. Furthermore, we developed a non-adversarial privacy protection strategy that prioritizes the preservation of semantic fidelity for legitimate receivers while constraining the semantic information accessible to adversaries. This approach, combined with an adaptive noise injection mechanism, enables the system to maintain a clear distinction between semantic and syntactic information recovery, thereby advancing the understanding of their respective roles in privacy and communication. Experimental results also highlight the limitations of traditional image quality metrics, such as PSNR and SSIM, in reflecting task-relevant semantic fidelity, highlighting the utility of our proposed metric. Overall, this work provides a unified framework that balances task optimization and privacy protection, contributing to the development of secure and efficient semantic communication systems.
Appendix A Appendix A

### v2

@@ IV-C Results on CIFAR-10 dataset
We also implement the system on the CIFAR-10 dataset, which has complex scenes that pose more challenges for semantic communication systems. Fig. 9 illustrates the receiver’s classification accuracy and the attacker’s success rate as functions of $t$. Given our assumption that the training of the $\mathcal{T}$-$\mathcal{R}$ pair takes precedence over the adversary’s training, even at $t=0$, the classification accuracy of the adversary’s reconstructed images is not very high. This is because the $\mathcal{T}$-$\mathcal{R}$ pair’s training prioritizes the completion of the label classification task, and the backpropagation process optimizes the $\mathcal{T}$’s neural network in a way that is not aligned with the adversary’s image recovery requirements, which is acceptable since normal systems are not designed to meet attackers’ needs.
@@
The two red lines in Fig. 9 show that the system incorporating the diffusion module outperforms the classifier-only scenario. This indicates that imperfect labels can still provide coarse-grained constraints for the diffusion path. However, the additional performance improvement brought by E-DiffSem is less pronounced compared to the MNIST dataset. This discrepancy may be because the CIFAR-10 dataset contains richer and more complex information than MNIST, which makes it more challenging for the imperfect labels estimated by the classifier to provide effective guidance to the diffusion module. In the future, we could explore integrating Transformer-based self-attention mechanisms or leveraging image semantic segmentation techniques to provide stronger guidance for the diffusion module, thereby further enhancing system performance.
@@
Figure 10: Plot of accuracy versus SSIM and PSNR using CIFAR-10 dataset.
Figure 11: Adversary’s output images with different $t$ under CIFAR-10 dataset. (a) Original image. (b)(c)(d) Adversary’s recovery when $t=60$, $150$ and $300$.
Fig. 10 shows the accuracy versus SSIM and PSNR on the CIFAR-10 dataset. It highlights the weak correlation between traditional image quality metrics and semantic security. For instance, when SSIM is $0.12$, the attacker’s classification accuracy still reaches $51.8\%$, whereas in the direct mode, with SSIM equal to $0.14$, the accuracy is only $42.6\%$. This difference arises from the attacker’s ability to implicitly learn class-discriminative features, while variations in pixel texture are insufficient to prevent the attacker from acquiring these features. Other curves in this figure can further demonstrate that SSIM and PSNR are not suitable metrics for evaluating the amount of task-relevant information obtained by the attacker in task-oriented communication systems.
As a supplement to the discussion, we comment on the computational complexity of our proposed system. Our hardware conditions are: CPU, Intel Xeon Gold 5220; GPU, GeForce RTX 3090. Although the diffusion model introduces additional computational overhead, the DDIM acceleration strategy restricts the additional single-sample inference time of DiffSem to approximately $0.1$s per batch (batch size 256) on CIFAR-10. It shows that the performance enhancement does not impose a substantial computational burden given the relatively small dataset. For higher-resolution datasets, such as ImageNet, the trade-off between performance improvement and computational cost may become a critical consideration, especially in real-time applications where latency is a key factor. Future work could focus on optimizing the diffusion process, potentially exploring techniques such as knowledge distillation, model pruning, or adopting more efficient neural network architectures to address this challenge.
@@ V Conclusion
In this paper, we introduce DiffSem, a diffusion-aided framework for task-oriented semantic communication, designed to balance task performance and privacy preservation. We generalize the algorithm by defining the success rate of model inversion via the attacker’s task accuracy, rather than pixel-level similarity metrics. DiffSem instantiates a non-adversarial privacy preserving mechanism that exploits the inherent mismatch between the task-specific semantics required by the legitimate receiver and the broader context that the attacker must reconstruct. Specifically, a transmitter-side self-noise mechanism adaptively regulates the radiated semantic information while compensating for channel noise, an unguided diffusion U-Net on the receiver enhances performance on downstream tasks, and an optional self-referenced label embedding further enhances receiver performance while maintaining the guidance-free constraint. Experiments show that DiffSem enables the legitimate receiver to achieve higher accuracy than the baseline, and highlight the potential mismatch between traditional image quality metrics and task-oriented semantic fidelity. Overall, DiffSem provides a unified and non-adversarial approach to balance utility and privacy in semantic communication, while our attacker accuracy-based evaluation method provides a more appropriate perspective for evaluating and mitigating model reversal risks in practice.
Appendix A Derivation of Eq. (24)